OpenStudy (anonymous):

Find the exact value of the inverse trigonometric function cos(sin^-1(root 2/3))

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Let \[\Large \theta = \sin^{-1}\left(\frac{\sqrt{2}}{3}\right)\] So that means \[\Large \sin(\theta) = \frac{\sqrt{2}}{3}\]

jimthompson5910 (jim_thompson5910):

since \[\Large \sin(\theta) = \frac{\sqrt{2}}{3}\] this means we can draw out this triangle 

jimthompson5910 (jim_thompson5910):

what is the missing side?

OpenStudy (anonymous):

wait how did you know what value corresponds with what angle/side?

jimthompson5910 (jim_thompson5910):

because sine = opposite/hypotenuse